# Status Report on the Chicago-Carnegie Hubble Program (CCHP): Three Independent Astrophysical Determinations of the Hubble Constant Using the James Webb Space Telescope

### v1

@@ 8 Comparison of the Cepheid, TRGB and JAGB Distances
In Table 2, we list the supernovae, host galaxy name, individual distances plus their uncertainties for each of the methods. Detailed descriptions of the measurement of these distances are provided in O24, H24 and L24. Here we compare the three measurements. We note that for the galaxy NGC 5643, there are two SNe Ia that have been observed, SN 2013aa and SN 2017cbv. For the Cepheid and TRGB calibrations, there are 11 SN Ia calibrators that have been observed as part of Pantheon+, and 10 for the CSP. (The CSP program ended in 2015 before SN 2017cbv was discovered.) For the JAGB calibration there are ten SN Ia calibrators, but three of these are of significantly lower weight. As described in L24, for three of the galaxies in our sample (NGC 3972, NGC 4038, and NGC 4424), no convergence in the radial measurement of the luminosity function was found.
@@
An immediate result from our comparisons of the three methods is that the TRGB and JAGB distances are in superb accord, whereas the measured Cepheid distances are offset (closer) than the other two. The upper panel of Figure 7 compares JAGB and TRGB distance moduli in our sample. The weighted (unweighted) mean difference between the JAGB minus TRGB distance moduli is 0.017 0.031 (0.013 0.013) mag or 1%. The scatter about the unit slope is 0.031 mag (1.4%). Immediately below, the residuals from the unit slope line are shown.
In the middle panel of Figure 7 we compare the Cepheid and TRGB distance moduli in our sample. In this case, the agreement is not as good, but is still reasonable: the scatter about the unit slope is 0.083 mag (3.9%). The weighted (and unweighted) mean difference between the Cepheid minus TRGB distance moduli is -0.054 0.028 (-0.049 0.028) mag, or a difference of about 2.5%.
Finally, the lower panel of Figure 7 compares the individual JAGB and Cepheid distance moduli in our sample. The scatter about the unit slope is 0.070 mag (3.3%). The weighted (unweighted) mean difference between the JAGB minus Cepheid distance moduli is 0.086 0.028 (0.083 0.031) mag or 4%.
We note that our Cepheid distances are in good agreement with those measured by R22. In Figure 8, we show a comparison of Cepheid distances measured in this study with those of R22. In this case, the weighted mean offset is -0.022 mag, amounting to a difference in distance (and ultimately ) of only 1%. The difference is in the sense of the new JWST distances being smaller.
The agreement for two of the (independent) methods (the JAGB and TRGB distances) is encouraging. The implication of these comparisons is that the TRGB and JAGB calibrations then lead to a lower value of than for the Cepheids; i.e., there is a local distance scale tension, independent of the tension in cosmology. These methods are all being applied to the same galaxies, for which there is a single distance. Understanding the reasons for these offsets is essential before we will be able to confirm if the Hubble tension is real. Importantly, we note that these differences arise in the distances themselves, before any application to the SNe Ia distance scale or determination of .
In concluding the discussion in this section, we note that there are several improvements to the analysis of Cepheid distances obtained as part of this study over those measured with HST alone, but in addition, there is a disadvantage. The advantages include the fact that (1) the new JWST near-infrared data have higher S/N (by more than an order of magnitude) than the previous HST near-infrared data. (2) The resolution is increased by a factor of four. (3) The re-analysis of the optical HST data has the benefit of the input source catalog coming from JWST. Hence, more stars contaminating the Cepheids can be removed in these new reductions. (4) As described in more detail in O24, and illustrated in Figure 9, three bands (F555W, F814W and F115W) provide a well-delineated extinction curve from which a correction for dust and a (total) reddening-corrected distance modulus can be obtained (e.g., Freedman, 1988; Freedman et al., 1991; Monson et al., 2012; Scowcroft et al., 2013). The dust corrections are consistent with what is expected for the foreground reddenings to these target galaxies, plus an additional contribution from dust internal to the target galaxy. However, a major disadvantage of the current JWST data is the smaller number of Cepheids observed in the calibrating galaxy, NGC 4258. The sampling within the Leavitt relation is such that a classical (least-squares) fit to the PL is very sensitive to the choice of short-period cut to the data. The uncertainty in the Cepheid distances for this calibration are thus larger than those in R22.
In §9 we provide a determination of based on each of the three distance scales individually, as well as a combined value based on the three methods.
@@ 9 Calibration of SNe Ia and the Hubble Constant
None of the three methods considered here (Cepheids, TRGB or JAGB stars) are bright enough to determine distances out into the smooth cosmic Hubble flow at a level of 1% accuracy in . Nearby galaxies and clusters induce motions or peculiar velocities, scattering above and below the Hubble expansion velocities, adding noise and potential bias to the determination of . In the past couple of decades, SNe Ia have surfaced as the preferred secondary distance indicator given their high intrinsic brightness and their small observed dispersion in the Hubble diagram (0.1-0.15 mag, e.g., (Burns et al., 2018; Scolnic et al., 2022; Brout et al., 2022)). We apply our independent Cepheid, TRGB and JAGB star distances to two samples of distant supernovae: 1) the ( CSP) (Uddin et al., 2024) and 2) the Pantheon+ sample (Scolnic et al., 2022), as described below.
@@ 9.1 The Carnegie Supernova Program (CSP)
The CSP was initiated 20 years ago as a program to provide multi-wavelength follow-up observations of previously discovered supernovae. The data were intended for applications to cosmology, as well as for studying the physical properties of the supernovae themselves666The CSP data are available at http://csp.obs.carnegiescience.edu/data.. The aim was to obtain homogeneous, intensive, high-cadence, multi- observations (Contreras et al., 2010). Careful attention was given to photometric precision and systematics: the program utilized a fixed set of instruments, photometric standard stars, and instrumental reduction procedures, catching most of the supernovae well before maximum, and with high signal to noise, avoiding many of the challenges otherwise faced in minimizing systematic differences between multiple data sets/instruments/etc. (Krisciunas et al., 2017). Optical spectra were also obtained with high cadence (Folatelli et al., 2013; Morrell et al., 2024). The bulk of the observations were carried out at Las Campanas Observatory using the 1-m Swope and 2.5-m du Pont telescopes. The first part of the CSP ( CSP-I) was carried out from 2004-2009, with a second phase ( CSP-II) from 2011-2015, optimized for the near-infrared (Phillips et al., 2019; Hsiao et al., 2019). Light curves were generated using the analysis package SNooPy (Burns et al., 2018). This program determines the peak magnitude in the light curves for each of the filters, the times of those maxima, (B-V) colors, and a color-stretch parameter sBV, as described by (Burns et al., 2014). The color-stretch parameter incorporates color in the relation between the luminosity and decline rate of a SN Ia, allowing also for the redder colors of fast decliners. Finally it computes all of the covariances amongst the parameters.
Previous applications of the CSP-I survey to cosmology include Burns et al. (2018), who used the CSP-I SNe Ia sample, calibrated using the Cepheid distances of Riess et al. (2016), to obtain a value of = 73.2 2.3 km s-1 Mpc-1 (for H-band data); and a value of = 72.7 2.1 km s-1 Mpc-1 (for B-band data). Calibrating the CSP-I sample with the TRGB (Freedman et al., 2019), updated in (Freedman et al., 2020; Freedman, 2021) found a lower value of = 69.8 0.6 (stat) 1.6 (sys) km s-1 Mpc-1. Uddin et al. (2024) have updated the CSP analysis, including the more recent CSP-II SNe Ia data, which, once included with CSP-I, triples the sample size over CSP-I. They undertook a calibration based on Cepheids, the TRGB and Surface Brightness Fluctuations, the latter a secondary distance indicator calibrated first by Cepheids. Using B-band light-curve fits, they find = 73.38 0.73 km s-1 Mpc-1 based on the Cepheid calibration of Riess et al. (2022). For the TRGB calibration, they find = 69.88 0.76 km s-1 Mpc-1, based on the TRGB distances of Freedman et al. (2019). Both of these results are in excellent agreement with the original studies.
@@ 9.2 JWST CSP calibration
The SNe Ia data used in this work are those described in the recent analysis by Uddin et al. (2024). There are over 300 SNe Ia in this sample including both CSP-I and CSP-II at wavelengths, and more than 200 SNe Ia observed at . In analyzing these data, we follow the methodology described previously by Burns et al. (2018); Freedman et al. (2019); Uddin et al. (2024). As in Burns et al., the corrected magnitudes for an individual filter (e.g., the B filter below) are given by:
@@
where in this case = , the apparent peak (K-corrected) magnitude, is the linear coefficient and is a quadratic coefficient in (); is the slope of the color correction; is the apparent peak magnitude at , K-corrected; and is the slope of the correlation between peak luminosity and host stellar mass . Host stellar masses are derived as described in Uddin et al. (2024). Following Uddin et al., the sample is split at the median mass so that equal weights are given above and below the median mass (although the results are not significantly affected by the choice of split point).
@@
The apparent magnitudes at maximum are computed by fitting the light-curves with SNooPy, providing the time of maximum, the light-curve shape , and the magnitude at maximum for each filter. These quantities are then provided as inputs to a Markov Chain Monte Carlo (MCMC) sampler that simultaneously solves for all the correction factors: , , , and (for full details, see Burns et al., 2018; Uddin et al., 2024). The MCMC sampler then provides the corrected magnitudes, as well as a full covariance matrix, which is used when determining and its error.
The distance modulus is defined as in (Burns et al., 2018) for a flat cosmology:
where is the deceleration parameter.
@@
As described in Uddin et al. (2024), three error terms are included in the analysis. The first, is the sum of the individual errors for the observed quantities, in addition to the covariance between the peak magnitude, and both the color and the color-stretch parameter. The second, is the intrinsic random scatter that allows for variations beyond either or those due to peculiar velocities, . The third term is the error due to the uncertainty in the distance measurements resulting from galaxy peculiar velocities, and scales with redshift. As in Burns et al. (2018), this term is given by where Vpec is a free parameter that represents the average peculiar velocity of the SN Ia sample.
@@
There are some differences in the CSP-I and CSP-II samples previously noted by Uddin et al. (2024). The scatter in the CSP-II data is larger than for CSP-I, which as noted by Uddin et al., results from a doubling of the number of low-redshift SNe Ia in the combined CSP sample. The average peculiar velocity (vpec) for the CSP-II sample is also larger than found for CSP-I, a consequence of increasing the number of low-redshift SNe Ia in the combined sample. The CSP-II sample doubled the total number of low-redshift SNe Ia with z (with 0.01 (14 SNe Ia in CSP-I, and 13 SNe Ia in CSP-II). To avoid issues with the largest peculiar velocities for nearby objects, in this paper, we limit the redshift to z 0.01 (e.g., see Brout et al., 2022). The parameter is also steeper in the combined sample than the CSP-I sample. One possibility is that the combined sample could contain more star-forming galaxies with larger amounts of dust. These issues highlight the fact that in the future, in an era where a value of at a level of 1% accuracy is the goal, large and homogeneous samples of SNe Ia become imperative.

### v2

@@ 8 Comparison of the TRGB and JAGB Distances
Here we compare the individual JWST JAGB and TRGB measurements. We note that for the galaxy NGC 5643, there are two SNe Ia that have been observed, SN 2013aa and SN 2017cbv. For the TRGB calibration, there are 11 SN Ia calibrators that have been observed as part of Pantheon+ and the CSP. For the JAGB calibration three of these are of significantly lower weight: as described in L25, for NGC 3972, NGC 4038, and NGC 4424, no convergence in the radial measurement of the luminosity function was found.
@@
An immediate result from our comparison is that the TRGB and JAGB distances are in superb agreement. This is demonstrated in Figure 7 that compares the eight overlapping JAGB and TRGB distance moduli in our sample (see Table 2). The weighted (unweighted) mean difference between the JAGB minus TRGB distance moduli is -0.003 0.019 (-0.008 0.018) mag or 1%. The scatter about the unit slope is 0.046 mag (2.1%). In the lower subpanel the residuals from the unit slope line are shown. The agreement for the two independent methods (the JAGB and TRGB distances) is encouraging.
9 Combining Previously-Published Observations of TRGB galaxies: I. HST Data
In this section, we describe, compare and combine the TRGB distances from F19 with our new JWST distances. This doubles our TRGB calibration sample from 10 to 20 galaxies, hosting a total of 24 SNe Ia, increasing the statistical precision of our TRGB determination. F19 utilized HST/ACS F814W (-band) observations to measure 18 SN Ia calibrators anchored to the Large Magellanic Cloud (LMC). F21 expanded the F19 sample with two new galaxies from Hoyt et al. (2021), excluding SN 2007on, underluminous by 3 relative to the calibrator sample, resulting in 19 SN Ia calibrators. Furthermore, F21 incorporated the Milky Way, SMC and NGC 4258 as additional geometric anchors to the calibration, increasing the total to four. Notably, both F19 and F21 share a common TRGB absolute magnitude zero point, M814W= -4.049 mag. Importantly, the four-anchor calibration agrees with the NGC 4258 zero point to within 0.001 mag.
Later in §14.2, we compare the TRGB distances for six galaxies in common to the I-band HST study and our current JWST -band study. Here we anticipate the excellent agreement at the 0.02 mag (1%) level. The absence of a significant systematic difference between the F19+F21 HST and our current H25 JWST calibrations of the TRGB allows us to combine our HST and JWST samples.
H25 further compared our -band TRGB distances with those of Anand et al. (2022) (the Extragalactic Distance Database or EDD), also finding good agreement with a weighted average offset difference of only 0.017 mag, albeit with larger scatter. The Anand et al. study is based on the same archival HST data, but the observations have been reduced independently using different software packages. This independent external check provides additional justification for combining our HST and JWST TRGB samples.
Hence, both of our HST and JWST distances are consistently tied to a common zero point set by the geometric parallax distance to NGC 4258, which is also in agreement with the more extensive calibration of the HST distances based on four geometric anchors. We list the total sample of CCHP TRGB distances in Table 3. In the cases where more than one TRGB measurement was made per galaxy, we took the inverse weighted average and weighted standard deviation (as described in §8) for our analysis.
| SN | Galaxy | ||||||||
| 2011fe | M101 | 29.151 | 0.042 | 29.078 | 0.04 | 29.113 | 0.029 | 29.194 | 0.039 |
| 2002fk | N1309 | … | … | 32.499 | 0.07 | 32.499 | 0.070 | 32.546 | 0.060 |
| 2006dd | N1316 | … | … | 31.460 | 0.04 | 31.460 | 0.040 | … | … |
| 1980N | N1316 | … | … | 31.460 | 0.04 | 31.460 | 0.040 | … | … |
| 1981D | N1316 | … | … | 31.460 | 0.04 | 31.460 | 0.040 | … | … |
| 2012fr | N1365 | 31.366 | 0.069 | 31.360 | 0.05 | 31.362 | 0.040 | 31.379 | 0.057 |
| 2011iv | N1404 | … | … | 31.360 | 0.06 | 31.360 | 0.060 | … | … |
| 2007on | N1404 | … | … | 31.360 | 0.06 | 31.360 | 0.060 | … | … |
| 2001el | N1448 | … | … | 31.320 | 0.06 | 31.321aaTRGB distances were obtained for four galaxies (NGC 1448, NGC 1559, NGC 5584 and NGC 5643) from F090W archival data discussed in §7.2 (not shown as columns). | 0.038 | 31.290 | 0.037 |
| 2021pit | N1448 | … | … | … | … | 31.321aaTRGB distances were obtained for four galaxies (NGC 1448, NGC 1559, NGC 5584 and NGC 5643) from F090W archival data discussed in §7.2 (not shown as columns). | 0.038 | 31.290 | 0.037 |
| 2015F | N2442 | 31.646 | 0.097 | … | … | 31.646 | 0.097 | 31.457 | 0.065 |
| 1995al | N3021 | … | … | 32.221 | 0.05 | 32.221 | 0.050 | 32.475 | 0.160 |
| 1998bu | N3368 | … | … | 30.313 | 0.04 | 30.313 | 0.040 | … | … |
| 1994ae | N3370 | … | … | 32.273 | 0.05 | 32.273 | 0.050 | 32.123 | 0.052 |
| 1989B | N3627 | … | … | 30.221 | 0.04 | 30.221 | 0.040 | … | … |
| 2011by | N3972 | 31.747 | 0.068 | … | … | 31.747 | 0.068 | 31.644 | 0.090 |
| 2007sr | N4038 | 31.645 | 0.078 | 31.681 | 0.05 | 31.671 | 0.042 | 31.615 | 0.117 |
| 2012cg | N4424 | 30.926 | 0.030 | 31.000 | 0.06 | 30.941 | 0.027 | 30.856 | 0.130 |
| 1994D | N4526 | … | … | 30.998 | 0.07 | 30.998 | 0.070 | … | … |
| 1981B | N4536 | 30.923 | 0.052 | 30.964 | 0.05 | 30.944 | 0.036 | 30.838 | 0.051 |
| 1990N | N4639 | 31.774 | 0.073 | … | … | 31.774 | 0.073 | 31.818 | 0.085 |
| 2007af | N5584 | … | … | 31.822 | 0.10 | 31.845aaTRGB distances were obtained for four galaxies (NGC 1448, NGC 1559, NGC 5584 and NGC 5643) from F090W archival data discussed in §7.2 (not shown as columns). | 0.047 | 31.775 | 0.053 |
| 2013aa | N5643 | 30.643 | 0.071 | 30.475 | 0.08 | 30.583aaTRGB distances were obtained for four galaxies (NGC 1448, NGC 1559, NGC 5584 and NGC 5643) from F090W archival data discussed in §7.2 (not shown as columns). | 0.039 | 30.570 | 0.050 |
| 2017cbv | N5643 | 30.643 | 0.071 | 30.475 | 0.08 | 30.583aaTRGB distances were obtained for four galaxies (NGC 1448, NGC 1559, NGC 5584 and NGC 5643) from F090W archival data discussed in §7.2 (not shown as columns). | 0.039 | 30.570 | 0.050 |
| 2013dy | N7250 | 31.629 | 0.047 | … | … | 31.629 | 0.047 | 31.628 | 0.126 |
9.1 Comparison of the Extended TRGB Sample with Riess et al. (2022)
Figure 8 compares our combined HST plus JWST TRGB distances with the R22 Cepheid distances for 17 common galaxies. The agreement is excellent, with a weighted mean difference of only 0.025 0.021 mag, corresponding to a 1.2% difference in distance. The scatter is 0.08 mag, or 4% in distance. The offset, at 1.2-, is consistent with zero. Both the CCHP TRGB and R22 Cepheid distance moduli for these galaxies are listed in Table 3.
This agreement in nearby distances is a noteworthy development. Previously, F19 reported a weighted average distance modulus difference (TRGB minus Cepheid) of +0.059 mag between the TRGB and Riess et al. (2019) distances. F19 noted also that the scatter in the local Hubble diagram was larger for the Cepheid distances, indicating that the discrepancy was not equally distributed between the two methods, suggesting that the TRGB distances had higher precision than the Cepheid distances. This 3% distance difference partially accounted for the discrepancy at the time ( = 73.24 1.74 and = 69.8 0.8 (stat) 1.7 (sys), for Cepheids and the TRGB, respectively). As outlined in F19, the differences in stemmed from: 1) the differences in distance moduli between the TRGB and Cepheid distances, 2) differing SNe Ia calibrators, and 3) varying weights assigned to SN Ia calibrators, arising from different uncertainties in distance and SN Ia measurements.
The first reason, the 0.06 mag discrepancy between the F19 TRGB and the R16 Cepheid distance scales, is now (partially) resolved as a source of the disagreement. As we have seen, the re-analysis by R22 of the Cepheid distances now results in distances to the SNe Ia that agree to within 0.025 mag of those from both the earlier F19 study and from our new JWST CCHP data (L25, H25).
@@ 10 Calibration of SNe Ia and the Hubble Constant
None of the three methods in our JWST program (Cepheids, TRGB or JAGB stars) are bright enough to determine distances out into the smooth cosmic Hubble flow at a level of 1% accuracy in . Nearby galaxies and clusters induce motions or peculiar velocities, scattering above and below the Hubble expansion velocities, adding noise and potential bias to the determination of . In the past couple of decades, SNe Ia have surfaced as the preferred secondary distance indicator given their high intrinsic brightness and their small observed dispersion in the Hubble diagram (0.1-0.15 mag, e.g., (Burns et al., 2018; Scolnic et al., 2022; Brout et al., 2022)). In this paper, we apply our independent TRGB and JAGB star distances to two samples of distant supernovae: 1) the ( CSP) (Uddin et al., 2024) and 2) the Pantheon+ sample (Scolnic et al., 2022), as described below.
@@ 10.1 The Carnegie Supernova Program (CSP)
The CSP (Contreras et al., 2010; Folatelli et al., 2010; Krisciunas et al., 2017) was initiated 20 years ago as a program to provide multi-wavelength follow-up observations of previously discovered supernovae. The data were intended for applications to cosmology, as well as for studying the physical properties of the supernovae themselves777The CSP data are available at http://csp.obs.carnegiescience.edu/data.. The aim was to obtain homogeneous, intensive, high-cadence, multi-band observations (Contreras et al., 2010). Careful attention was given to photometric precision and systematics: the program utilized a fixed set of instruments, photometric standard stars, and instrumental reduction procedures, catching most of the supernovae well before maximum, and with high signal-to-noise, avoiding many of the challenges otherwise faced in minimizing systematic differences across multiple data sets/instruments/etc. (Krisciunas et al., 2017). Optical spectra were also obtained with high cadence (Folatelli et al., 2013; Morrell et al., 2024). The bulk of the observations were carried out at Las Campanas Observatory using the 1-m Swope and 2.5-m du Pont telescopes. The first part of the CSP ( CSP-I) was carried out from 2004-2009, with a second phase ( CSP-II) from 2011-2015, optimized for the near-infrared (Phillips et al., 2019; Hsiao et al., 2019). Light curves were generated using the analysis package SNooPy (Burns et al., 2018). This program determines the peak magnitude in the light curves for each of the filters, the times of those maxima, (B-V) colors, and a color-stretch parameter sBV, as described by Burns et al. (2014). The color-stretch parameter incorporates color in the relation between the luminosity and decline rate of a SN Ia, allowing also for the redder colors of fast decliners. Finally, it computes the covariance matrix involving all of the parameters.
Previous applications of the CSP-I survey to cosmology include Burns et al. (2018), who used the CSP-I SNe Ia sample, calibrated using the Cepheid distances of Riess et al. (2016), to obtain a value of = 73.2 2.3 km s-1 Mpc-1 (for H-band data); and a value of = 72.7 2.1 km s-1 Mpc-1 (for B-band data). Calibrating the CSP-I sample with the TRGB (Freedman et al., 2019), updated in (Freedman et al., 2020; Freedman, 2021) found a lower value of = 69.8 0.6 (stat) 1.6 (sys) km s-1 Mpc-1. Uddin et al. (2024) have updated the CSP analysis, including the more recent CSP-II SNe Ia data, which, once included with CSP-I, triples the sample size over CSP-I. They undertook a calibration based on Cepheids, the TRGB and Surface Brightness Fluctuations, the latter being a secondary distance indicator calibrated first by Cepheids. Using B-band light-curve fits, they find = 73.38 0.73 km s-1 Mpc-1 based on the Cepheid calibration of Riess et al. (2022). For the TRGB calibration, they find = 69.88 0.76 km s-1 Mpc-1, based on the TRGB distances of Freedman et al. (2019). Both of these later results are in excellent agreement with the original studies.
@@ 10.2 JWST CSP calibration
The SNe Ia data used in this work are those described in the recent analysis by Uddin et al. (2024). There are over 300 SNe Ia in this sample, including both CSP-I and CSP-II at wavelengths, and more than 200 SNe Ia observed at . In analyzing these data, we follow the methodology described previously by Burns et al. (2018); Freedman et al. (2019); Uddin et al. (2024). As in Burns et al., the corrected magnitudes for an individual filter (e.g., the B filter below) are given by:
@@
where in this case = , the apparent peak (K-corrected) magnitude, is the linear coefficient, and is a quadratic coefficient in () where is the stretch parameter described in §10.1; is the slope of the color correction; is the apparent peak magnitude at , K-corrected; is the slope of the correlation between peak luminosity and host stellar mass ; and M0 is the median value of the host stellar mass. Host stellar masses are derived as described in Uddin et al. (2024). Following Uddin et al., the sample is split at the median mass so that equal weights are given above and below the median mass (although the results are not significantly affected by the choice of split point).
@@
The apparent magnitudes at maximum are computed by fitting the light curves with SNooPy, providing the time of maximum, the light-curve shape , and the magnitude at maximum luminosity for each filter. These quantities are then provided as inputs to a Markov Chain Monte Carlo (MCMC) sampler that simultaneously solves for all the correction factors: , , , and (for full details, see Burns et al., 2018; Uddin et al., 2024). The MCMC sampler then provides the corrected magnitudes, as well as a full covariance matrix, which is used when determining and its error.
The distance modulus is defined as in Burns et al. (2018) for a flat cosmology:
where is the deceleration parameter, and z is the redshift in the CMB frame.
@@
As described in Uddin et al. (2024), three error terms are included in the analysis. The first, is the sum of the individual errors for the observed quantities, in addition to the covariance between the peak magnitude, and both the color and the color-stretch parameter. The second, is the intrinsic random scatter that allows for variations beyond either or those due to peculiar velocities, . The third term is the error due to the uncertainty in the distance measurements resulting from galaxy peculiar velocities, and scales with redshift. Burns et al. (2018) define where Vpec represents the average peculiar velocity of the SN Ia sample. We set the average peculiar velocity Vpec = 240 km s-1, a value consistent with that determined by Brout et al. (2022); Scolnic et al. (2015); Burns et al. (2018). In practice, this made an insignificant difference in the determination of . Adopting different values, taken in steps of 50 km s-1 in the range of 150 km s-1 resulted in differences in at a level 1%.
@@
There are some differences in the CSP-I and CSP-II samples previously noted by Uddin et al. (2024). The combined CSP-I plus CSP-II sample results in an increase by almost a factor of three relative to the CSP-I sample alone. The scatter in the CSP-II data is larger than for CSP-I, which as noted by Uddin et al., results from a doubling of the number of low-redshift SNe Ia in the combined CSP sample. The average peculiar velocity (Vpec) for the CSP-II sample was also larger than found for CSP-I. In this paper, we consider only redshifts of z 0.01 (see e.g., Brout et al., 2022; Chen et al., 2024) to limit effects of peculiar velocities for nearby objects. We explored redshift cuts ranging from 0.0065 z 0.03, and found differences with a full range of only 1% in . These differences were stochastic, and showed no trend with increasing redshift cuts. Finally, the parameter is also steeper in the combined sample than the CSP-I sample. One possibility is that the combined sample could contain more star-forming galaxies with larger amounts of dust. These issues highlight the fact that in the future, in an era where a value of at a level of 1% accuracy is the goal, large and homogeneous samples of SNe Ia become increasingly important.